# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

### v4

@@ 2.5.1 Attributes Construction
First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work. Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntosh et al. (2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens. Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al. (2023), transforming them into factors for our experiment. Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations. In the end, we obtained a list of 10 factors in Table 3, along with the source, definition, and rationale for each factor.
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. We aggregated the dataset from three repetitive experiments after filtering out code files with multiple security defects. We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment, and their factor values were ready for model fitting.
@@ 2.5.2 Regression Model Construction
Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 3. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 3, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package [^2] https://cran.r-project.org/web/packages/rms/index.html, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating Spearman’s rank correlations (${\rho}^{2}$) between them. As  Harrell (2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned higher degrees of freedom.  Paul et al. (2021) suggested limiting the maximum allocated degrees of freedom for an explanatory variable to five or fewer, thus reducing the risk of overfitting. According to the computation results shown in Fig. 4, for interval variables, we considered using the rcs function from the R’s rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
@@
Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015). We tested models with and without allocating 3 degrees of freedom to the ‘Token’ factor to balance model complexity and fit. The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2025).
@@ 3.1 Performance of LLMs (RQ1)
Based on the evaluation metrics defined in Section 2.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts. We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively. If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperformed the others. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 4 and Table 5), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 6). The best values for each score are highlighted in bold and marked with asterisks in the tables.
From Table 4 and Table 5, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 6 demonstrated that in the selected 57 cases, Llama 2 70B scored lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in the complete datasets. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools. This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions. As such, features can only be accurately modelled using dynamic analysis, and these security defects are challenging for static analysis tools to detect (Charoenwet et al., 2024).
@@
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
@@
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
@@
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
@@
M-Score 70.18% 80.12% 42.69% 20.47*44.44%
@@
When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4 and Table 5), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%). Following closely behind is Llama 2 70B and GPT-3.5. GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit. Similar results are observed in Table 6. Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, GPT-4 achieved I-Score = 8.77%, IH-Score = 20.46%, and M-Score = 43.86%. It is evident that GPT-4’s performance on the selected 57 instances exceeds its performance across the entire dataset. Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
@@
As shown in Table 4 and Table 5, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small. Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively. Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement. A possible reason is that the name of components or packages utilized may already imply the source project of the code file, thereby diminishing the impact of project information provided in $\textbf{P}_{r}$. Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific. We speculate that some LLMs may already understand the task described in the prompt quite well, and thus, introducing a persona could add unnecessary complexity to the prompt and lead to distractions.
We can see from Table 4 and Table 5 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflects that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects. Still, it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, leveraging CWE information in the prompt is worthy of consideration.
@@
Table 4 and Table 5 illustrate that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B in both Python and C/C++ datasets. However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs. As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets. It can be seen that the impact of supplementing missing context by CoT prompting on the performance of LLMs is multifaceted and unstable. A possible explanation is that different LLMs vary in their ability to process and understand input. Moreover, we instructed the LLM to generate missing code context based on the corresponding commit message in $\textbf{P}_{cot}$. On the one hand, the generated context may contain the critical information needed to analyze the target security defect, which can help the LLM in security code review. On the other hand, the consistency between this generated context and the actual code context cannot be guaranteed. We observed that in some cases, the code context generated under Prompt $\textbf{P}_{cot-1}$ was inconsistent with the actual code context, resulting in the original security defect no longer being identified as a defect within the generated context. Consequently, when Prompt $\textbf{P}_{cot-2}$ was applied, the detection of security defects in the provided code was misled by the generated code context and yielded erroneous results.
@@
According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater non-determinism.

### v5

@@ 2.5.1 Attributes Construction
First, we collected the attributes from previous works (paul2021security; mcintosh2016empirical; thongtanunam2017review) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work. Given that LLMs understand and generate text token by token, we adjusted the factor adopted from mcintosh2016empirical — ‘Size’, to ‘Token’ (see the first row of Table 2), as larger code files typically contain more tokens. Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in chen2023chatgpt; Sovrano_2025, transforming them into factors for our experiment. Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations. In the end, we obtained a list of 11 factors in Table 2, along with the source, definition, and rationale for each factor.
[…]
@@ 2.5.2 Regression Model Construction
Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (baak2020new), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 6. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (hinkle2003applied). From Fig. 6, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (legendre2012numerical). Utilizing the redun function in the rms package [^2] https://cran.r-project.org/web/packages/rms/index.html, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating Spearman’s rank correlations (${\rho}^{2}$) between them. As  harrell2001regression recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned higher degrees of freedom.  paul2021security suggested limiting the maximum allocated degrees of freedom for an explanatory variable to five or fewer, thus reducing the risk of overfitting. According to the computation results shown in Fig. 7, for interval variables, we considered using the rcs function from the R’s rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
@@
[…]
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (replpack).
@@ 3.1 Performance of LLMs (RQ1)
Based on the evaluation metrics defined in Section 2.3.3, we evaluated the capability of six popular LLMs in security code review across different prompts. We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively. If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperformed the others. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To avoid bias due to different samples and to ensure rigorous comparisons, we also performed analysis on the intersection of source code files that all LLM-prompt combinations could generate responses, which yielded 57 instances. As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 3 and Table 4), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 5). The best values for each score are highlighted in bold and marked with asterisks in the tables. It is evident that, DeepSeek-R1 outperformed all other models, followed by GPT-4. Specifically, Table 3 shows that on the Python dataset, the combination of DeepSeek-R1 with $\textbf{P}_{cid}$ achieved the best performance. According to Table 4, on the C/C++ dataset, the combination of DeepSeek-R1 with $\textbf{P}_{cot-guardrail}$ performed the best. Furthermore, Table 5 demonstrates that in the selected 57 cases, DeepSeek-R1 with $\textbf{P}_{b}$ achieved better results than all other LLM-prompt combinations. Although GPT-4 did not perform as well as DeepSeek-R1, it still obviously surpassed the remaining five LLMs. For the top two models, DeepSeek-R1 and GPT-4, we calculated the performance scores of these two models across different prompt templates on the full dataset (including both Python and C/C++ datasets), as shown in Table 6. We found that DeepSeek-R1 performed best with $\textbf{P}_{cot-guardrail}$, whereas GPT-4 achieved its highest scores with $\textbf{P}_{cid}$. On the whole, as of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools. This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (russo2010dynamic); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions. As such, features can only be accurately modelled using dynamic analysis, and these security defects are challenging for static analysis tools to detect (charoenwet2024empirical).
@@
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot-guardrail}$
@@
DeepSeek-R1 Resp. Count 234 234 234 234 234 234
I-Score 5.41% 6.27% 4.70% 7.26%* 4.56% 5.84%
IH-Score 7.83% 7.55% 6.27% 9.26% 6.41% 8.69%
M-Score 36.04% 32.05% 29.34%* 30.77% 46.87% 35.69%
Entropy 0.5440 0.5273 0.5098 0.5444 0.5683 0.5526
@@
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot-guardrail}$
@@
DeepSeek-R1 Resp. Count 269 269 269 269 269 269
I-Score 12.64% 8.43% 9.91% 10.61% 9.79% 12.89%*
IH-Score 15.49%* 10.41% 11.15% 12.27% 12.64% 15.49%*
M-Score 29.99% 20.94%* 21.93% 23.05% 34.32% 27.39
Entropy 0.5411 0.4934 0.4884 0.5002 0.5654 0.5663
@@
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot}$
@@
M-Score 70.18% 80.12% 42.69% 20.47*% 44.44%
DeekSeek-R1 I-Score 16.96%* 16.37% 12.87% 16.37% 13.45% 12.28%
IH-Score 23.98%* 19.88% 15.79% 21.64% 18.71% 18.13%
M-Score 40.35% 37.43% 35.67% 30.99% 45.61% 32.75%
Table 6: Performance scores and entropy for GPT-4 and DeepSeek-R1 under different prompts on the complete dataset
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot-guardrail}$
GPT-4 Resp. Count 422 422 422 421 421
I-Score 4.58% 3.71% 3.63% 5.30%* 4.35%
IH-Score 8.53% 8.63% 8.92% 12.91%* 10.29%
M-Score 57.89% 62.88% 45.33% 41.09%* 51.07%
DeepSeek-R1 Resp. Count 503 503 503 503 503 503
I-Score 9.28% 7.43% 7.49% 9.05% 7.36% 9.61%*
IH-Score 11.93% 9.08% 8.88% 10.87% 9.74% 12.33%*
M-Score 32.80% 26.11% 25.38%* 26.64% 40.16% 31.25
@@
When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 3 and Table 4), DeepSeek-R1 performed the best, obtaining the highest I-Score (Python: 5.41%, C/C++: 12.64%), and the lowest M-Score (Python: 36.04%, C/C++: 29.99%). For the IH-Score, DeepSeek-R1 still achieved the highest value on the C/C++ dataset, reaching 15.49%, while on the Python dataset, GPT-4 outperformed DeepSeek-R1 (7.83%) with a score of 9.34%. GPT-4 was regarded as the second-best model, followed by Llama 2 70B and GPT-3.5. GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit. Similar results are observed in Table 5. Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, DeepSeek-R1 achieved I-Score = 16.96%, IH-Score = 23.98%, and M-Score = 40.35%. It is evident that DeepSeek-R1’s performance on the selected 57 instances exceeds its performance across the entire dataset. Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
@@
$\textbf{P}_{r}$: As shown in Table 3 and Table 4, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small. Specifically, DeepSeek-R1, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively. Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement. A possible reason is that the name of components or packages utilized may already imply the source project of the code file, thereby diminishing the impact of project information provided in $\textbf{P}_{r}$. Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific. We speculate that some LLMs may already understand the task described in the prompt quite well, and thus, introducing a persona could add unnecessary complexity to the prompt and lead to distractions.
$\textbf{P}_{c}$: We can see from Table 3 and Table 4 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of DeepSeek-R1 and GPT-3.5 significantly decreases under $\textbf{P}_{c}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets.
$\textbf{P}_{cid}$: Under $\textbf{P}_{cid}$, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B and Llama 2 70B all achieved their best performance on both the Python and C/C++ datasets, while GPT-3.5 experienced a significant drop in scores. However, the impact of $\textbf{P}_{cid}$ on the performance of DeepSeek-R1 was language-specific: while it led to a decline in C/C++, it enhanced performance in Python. It can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of general-purpose LLMs in security code review. Notably, under $\textbf{P}_{c}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflects that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects. Still, it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, incorporating specific CWE entries in the prompt is worthy of consideration. For DeepSeek-R1, we analyzed why the impact of $\textbf{P}_{cid}$ on performance is programming language-specific. As mentioned in Section 2.2, Race Condition is the most prevalent security defect type in the Python dataset. The CWE list provided in $\textbf{P}_{cid}$ includes CWE-435, which is highly relevant to Race Condition, thereby enhancing model performance. In contrast, the C/C++ dataset features a more balanced distribution of security defect types, such as Integer overflow, Resource Leak and Crash. We included solely high-level CWE entries in the CWE list and some of them, such as CWE-682 (Incorrect Calculation), do not exactly match certain security defect types like Integer Overflow. These CWE entries failed to provide sufficient guidance and even introduced interference. More fine-grained subtypes may provide better guidance for specific security defect types. Therefore, determining the optimal granularity of CWE as auxiliary information for different LLMs in security defect detection warrants further investigation in future work.
@@
Table 3 and Table 4 illustrate that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B in both Python and C/C++ datasets. However, for Gemini Pro, GPT-4 Turbo and DeepSeek-R1, there is no improvement and even a detrimental effect to these LLMs. As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets. It can be seen that the impact of supplementing missing context by CoT prompting on the performance of LLMs is multifaceted and unstable. A possible explanation is that different LLMs vary in their ability to process and understand input. Moreover, we instructed the LLM to generate missing code context based on the corresponding commit message in $\textbf{P}_{cot}$. On the one hand, the generated context may contain the critical information needed to analyze the target security defect, which can help the LLM in security code review. On the other hand, the consistency between this generated context and the actual code context cannot be guaranteed. We observed that in some cases, the code context generated under Prompt $\textbf{P}_{cot-1}$ was inconsistent with the actual code context, resulting in the original security defect no longer being identified as a defect within the generated context. Consequently, when Prompt $\textbf{P}_{cot-2}$ was applied, the detection of security defects in the provided code was misled by the generated code context and yielded erroneous results.
To isolate the impact of LLM-generated code context on model performance, we designed $\textbf{P}_{cot-guardrail}$ and conducted experiments under this prompt on the best-performing model, DeepSeek-R1. $\textbf{P}_{cot-guardrail}$ retains CoT guidance and commit message but does not include the fabricated code context. The results are also presented in Table 3, Table 4 and Table 5. We observed that, on both the Python and C/C++ datasets, $\textbf{P}_{cot-guardrail}$ consistently improved the performance of DeepSeek-R1 compared with $\textbf{P}_{b}$, whereas $\textbf{P}_{cot}$ led to a performance decline. This suggests that for DeepSeek-R1, LLM-generated code context introduced interference for security code review. However, in the selected 57 instances, DeepSeek-R1 performed worse with $\textbf{P}_{cot-guardrail}$ than with $\textbf{P}_{b}$. The 57 samples satisfy the token limits of all LLM-prompt combinations in our experiments, making them the samples with the fewest tokens in the dataset. Based on this, we hypothesize the following explanation: the commit message included in $\textbf{P}_{cot-guardrail}$ helps the model understand code semantics and the CoT instruction drives the model to perform step-by-step reasoning, making this prompt more suitable for assisting the model in analyzing security defects in complex and long code. For code files with fewer tokens, these additional guidance could be unnecessary or even introduce noise to the LLM.
Missing code context can impact the reliability of model outcomes (li2025everything), which remains a challenge when applying LLMs to real-world security code review. We attempted to generate code context using the LLM itself, but did not directly measure how often the LLM hallucinated irrelevant code and how such hallucinations affected model performance. This is because evaluating the consistency between LLM-generated code and the actual source code is challenging, as it requires a comprehensive understanding of the codebase and its historical commits. We plan to address this in future work. Some works have performed context augmentation by incorporating call graphs extracted through static analysis tools (luo2024fellmvp) or by utilizing caller-callee relationships within the code (wen2024vuleval; gao2023far; sun2024llm4vuln). Given this, we will also compare these approaches with the LLM-driven code context generation in assisting LLMs with security code reviews, and explore how to better leverage the capabilities of LLMs to obtain more effective and precise code context.
@@
Figure 8: Response Entropy of each LLM-prompt combinations on the Python dataset
Figure 9: Response Entropy of each LLM-prompt combinations on the C/C++ dataset
According to the mean entropies and their 95% confidence intervals (CI) of various LLM-prompt combinations in Fig. 8 and Fig. 9, we found that GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B, Llama 2 70B and DeepSeek-R1, with GPT-4 showing the least consistency.
Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with Gemini Pro, with a mean entropy of 0.0404 (95% CI: [0.02, 0.07]). On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195 (95% CI: [0.00, 0.05]). Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code. On the Python dataset, the highest average entropy of 0.9416 (95% CI: [0.89, 0.99]) was observed with the combination of GPT-4 and $\textbf{P}_{cot}$. On the C/C++ dataset, GPT-4 paired with $\textbf{P}_{cid}$ exhibited the highest average entropy of 0.7330, with a 95% CI from 0.67 to 0.79. Compared with other LLMs excluding DeepSeek-R1, GPT-4 not only demonstrated significantly superior performance, but also exhibited the highest uncertainty. In comparison with GPT-4, DeepSeek-R1 showed improved performance in the security code review task but produced more stable outputs. However, the average entropy of DeepSeek-R1 remained relatively high among all models evaluated in our experiments. As shown in Table 1, we kept the decoding parameters at their default settings for all models. Except for DeepSeek-R1, the temperature and top_p values of all LLMs fell within a similar range (0.9–1.0). Despite this, we observed substantial differences in response consistency across models. Furthermore, although DeepSeek-R1 was configured with a lower top_p value (0.7), its response consistency was still relatively poor among all models. This observation suggests that decoding parameters alone may have a limited effect on regulating output consistency, and the observed differences are more likely attributable to model-intrinsic randomness.